OpenStudy (anonymous):

CAL 2:Find indefinite integral using simple substitution: Integral of (cosx/(1+(sinx)^2))dx

OpenStudy (anonymous):

\[\int\limits cosx/(1+(sinx)^2)dx\]

OpenStudy (anonymous):

yes u=sin(x) du=cos(x)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}1/(u ^{2}+1)du\]

OpenStudy (anonymous):

it is arctan(u)

OpenStudy (anonymous):

and u = sin(x)

OpenStudy (anonymous):

don't forgot constant )))

OpenStudy (anonymous):

full answer is arctan(sin(x))+const

OpenStudy (anonymous):

that simple? should I use first as denominator to change to 1+sin^2x and use identity?

OpenStudy (anonymous):

as sin^2x=1-cos^2x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lululz this aint sinh and cosh

OpenStudy (anonymous):

its called partly integration

OpenStudy (anonymous):

sorry it was not arctan it must be cotan

OpenStudy (anonymous):

right answer is cot(sin(x))+const

OpenStudy (anonymous):

ok I was thinking:\[\int\limits cosx/1+(1-\cos^2x) \to change given and than go from there but I see, \it would be more complicated\]

OpenStudy (anonymous):

no don't this

OpenStudy (anonymous):

its not cot it can't be I think inverse tan was right

OpenStudy (anonymous):

no its cot=tan^-1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

julia, where are you from?

OpenStudy (anonymous):

if not secret =)

OpenStudy (anonymous):

:O) Miami

OpenStudy (anonymous):

and where do you study?